2. A market consists of two population segments, A and B. An individual in segment A has demand for your product q = 50 – p. An individual in segment B has demand for your product q = 120 – 2p. Segment A has 1000 people in it. Segment B has 1200 people in it. Total cost of producing q units is C = 5000 + 20q.

1) What is total market demand for your product?

2) Assume that you must charge the same price to both segments. What is the profit-maximizing price? What are your profits?

3) Imagine now that members of segment A all wear scarlet “A” on their shirts or blouses and that you can legally charge different prices to these people. What price do you charge to the scarlet “A” people? What price do you charge to those without the scarlet “A”? What are your profits now?
1
Expert's answer
2015-09-18T00:00:45-0400
2. A market consists of two population segments, A and B. An individual in segment A has demand for your product q = 50 – p. An individual in segment B has demand for your product q = 120 – 2p. Segment A has 1000 people in it. Segment B has 1200 people in it. Total cost of producing q units is C = 5000 + 20q.
1) The total market demand for your product is: Qd = qA + qB = 50 - p + 120 - 2p = 170 - 3p
2) If you must charge the same price to both segments, then the profit-maximizing price will be in the point, where P = MR = MC, MC = C' = 20, so P = 20. Your profits are: TP = TR - C = P*q - C = 20*2200 - 5000 - 20*2200 = -5000.
3) If you can legally charge different prices to these people, the price you can charge to the scarlet “A” people is higher, so your profits will be higher too.